A triangle ABC has a perimeter of 1.89m. AB + BC = 2AC. AB = 2BC.Find the length of the longest side.
Answer by greenestamps(13367)



AB = 2BC

let BC = x
then AB = 2x

AB+BC = 2AC

2x+x = 2AC
3x = 2AC
AC = 1.5x

The perimeter is 1.89m:

2x+x+1.5x = 1.89
4.5x = 1.89
9x = 3.78
x = 3.78/9 = 0.42

The side lengths are
AB = 2x = 0.84m
BC = x = 0.42m
AC = 1.5x = 0.63m

ANSWER: The longest side is 0.84m